OpenStudy (anonymous):

Two sets of equatic expressions are shown below in various forms: Line 1: x2 + 3x + 2 (x + 1)(x + 2) (x + 1.5)2 − 0.25 Line 2: x2 + 5x + 6 (x + 2)(x + 3) (x + 2.5)2 + 6.25 Which line contains three equivalent expressions?

OpenStudy (anonymous):

Please help

OpenStudy (mathmale):

What does "equivalent expression" mean to you? How do you recognize an "equivalent expression?"

OpenStudy (anonymous):

it means the expressions have the same answer

OpenStudy (anonymous):

(equivalent answers)

OpenStudy (mathmale):

Not quite. expressions don't have "answers." What have you done so far? What could you do to each expression to give it just one format that you can easily compare?

OpenStudy (anonymous):

factor them?

OpenStudy (mathmale):

"answer" isn't applicable here. It's "equivalence" that's important. Have you tried factoring any of the given expressions?

OpenStudy (anonymous):

yeah

OpenStudy (mathmale):

I'd like to see your results. Please share whatever you've done so far.

OpenStudy (anonymous):

first one is (x+1)(x+2)

OpenStudy (mathmale):

Simplify as much as possible, and factor: (x + 1.5)2 − 0.25

OpenStudy (anonymous):

its (x+2)(x+1)

OpenStudy (anonymous):

or (x + 1)(x + 2)

OpenStudy (mathmale):

Note: by (x+1.5)2 you probably meant\[(x+1)^{2}\]

OpenStudy (anonymous):

no i think your wrong on this 1

OpenStudy (mathmale):

Yes, I'm wrong. Show me all of your own work.

OpenStudy (anonymous):

if you factor (x + 1.5)2 − 0.25 it will give you (x + 1)(x + 2)

OpenStudy (mathmale):

Expand (x+1.5)^2.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait expand???

OpenStudy (mathmale):

That (x + 1.5) is a binomial. You are asked to find the square of this binomial. That means multiplying out (x+1.5)(x+1.5). That's what I meant by "expansion"

OpenStudy (mathmale):

Expand:\[(x+1.5)^2\]

OpenStudy (anonymous):

(x+1.5)(x+1.5) =(x+1.5)(x+1.5) =(x)(x)+(x)(1.5)+(1.5)(x)+(1.5)(1.5) =x2+1.5x+1.5x+2.25 =x2+3x+2.25

OpenStudy (anonymous):

there

OpenStudy (mathmale):

Good. Please, indicate exponentiation with the ^ symbol. (x+1.5)^2, not (x+1.5)2.

OpenStudy (anonymous):

k

OpenStudy (mathmale):

Now, subtract .25 from your result. What do you get?

OpenStudy (anonymous):

x2+3x+2.25−0.25 =x2+3x+2.25+−0.25 Combine Like Terms: =x2+3x+2.25+−0.25 =(x2)+(3x)+(2.25+−0.25) =x2+3x+2 Answer: =x2+3x+2

OpenStudy (mathmale):

Good. Now, factor that.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

(x+1)(x+2)

OpenStudy (anonymous):

just like my first answer

OpenStudy (mathmale):

Look at the given expressions, first row. Is your most recent result the same as the middle expression in the first row?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i know that

OpenStudy (mathmale):

great. So, we call the 2nd and 3rd expressions in the first row " ? "

OpenStudy (anonymous):

so line 1 is a "equivalent expression"

OpenStudy (anonymous):

now we need to find out about line 2

OpenStudy (anonymous):

Line 2: x2 + 5x + 6 (x + 2)(x + 3) (x + 2.5)2 + 6.25

OpenStudy (mathmale):

We haven't looked at the first expression in line 1 yet. but "equivalent" was properly applied here. Are all 3 expressions in Line 1 equivalent? yes or no

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

So, Line 1 could be the line y ou were looking for, with all 3 expressions "equivalent."

OpenStudy (anonymous):

all of then are (x+1)(x+2)

OpenStudy (anonymous):

but we need to find out if line 2 is or not

OpenStudy (mathmale):

Please re-write (x + 2.5)2 + 6.25 using the proper symbol for exponentiation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x2+5x+12.5

OpenStudy (mathmale):

??? Where is your " ^ " symbol, to express exponentiation? Where did that 12.5 come from?

OpenStudy (anonymous):

nvm ur no help

OpenStudy (mathmale):

I'm making sure you understand the concepts involved. Sorry that you don't want to take this to completion, especially since we're so close to finishing.

OpenStudy (mathmale):

Please re-write (x + 2.5)2 + 6.25 using ^ for exponentiation, and explain how you got 12.5.